# Complete transparency with three active-passive-coupled optical resonators

### v1

@@ Abstract
In summary, we have presented a scenario of genuinely reducing the susceptibility $\varepsilon_{T}$ of the structure in Fig. 1(a) to zero under the implementable conditions, so that the described phenomenon of complete transparency is totally irrelevant to the input field power. Apart from scanning the input frequency, we only require the tuning of two inter-cavity couplings for seeing the phenomenon: setting one inter-cavity coupling at $J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ and adjusting the other coupling $J_{1}$ (possibly together with the external coupling $\kappa_{ex}$) for a suitable transparency window size. Rather than being used to cancel the transmission loss, the optical gain plays the role of determining the point $J_{2,c}$ together with the field damping rate in a neighboring resonator, so any nonzero effective gain rate $\kappa_{2}$ can be used in principle. The significance of the scenario lies in two aspects. First, to the research field of induced transparency based on optical resonator [22, 23], it proves that the ultimate goals of complete transparency and free control on the output/input ratio can be realized without complex designs or particular nonlinear elements. Second, it will non-trivially extend the scope of non-Hermitian photonics based on optical gain. Optical gain in one of two coupled optical resonators provides a realization of PT symmetry [47, 48, 49, 50, 51, 52, 60, 69], and the study of the associated phenomenon of exceptional points has been developed into a direction of intensive research [70, 71, 72]. Coupling one more resonator will have an even more non-trivial phenomenon involving a special dark mode with its rich properties, and it presents a rare example of something unusual simply built on the common elements like active, passive resonators and their linear couplings.
@@ Appendix A Appendix
where the noise operators satisfy the relations $\langle\hat{\xi}(t)\hat{\xi}^{\dagger}(t^{\prime})\rangle=\delta(t-t^{\prime})$ and $\langle\hat{\zeta}^{\dagger}(t)\hat{\zeta}(t^{\prime})\rangle=0$. In some other literature the damping terms take the form $\kappa/2$ while the corresponding noise terms have the factor $\sqrt{\kappa}$. Particularly, we consider the gain saturation effect such that the effective gate rate with time takes the form
@@
In some applications, gain saturation effect leads to optical isolation of coupled optical resonators [53, 56, 60], modifies optomechanical oscillations of hybrid systems [55, 57], and can also influence macroscopic quantum properties of coupled resonators [54]. In our concerned system of three coupled optical resonators, the time-dependent net optical gain rate can affect the intracavity fields. The indicator for the system to evolve to a desired dark mode is whether the field amplitude in the first cavity will finally tend to $|\langle\hat{a}_{1}\rangle|=0$, where $\langle\hat{a}_{1}\rangle=A_{1}e^{-i\omega_{p}t}$. We let the system evolve under two conditions: (1) the resonant input for three identical resonance cavity frequencies; (2) $J_{2}=\sqrt{\kappa_{2,0}\kappa_{3}}$ for the chosen system parameters. Without a loss of the described features, we approximate with $\gamma_{2}=0$ in our model calculations because this loss can be compensated by a higher $\kappa_{2,0}$, so that the net gain rate keeps to be $\kappa_{2}(t)>0$. For a constant $\kappa_{2}=\kappa_{2,0}$, there will emerge an exact $|\langle\hat{a}_{1}\rangle|=0$ under these conditions, as we illustrate in the main text, but this result cannot be guaranteed with a variable $\kappa_{2}(t)$ in Eq. (S-3). Through the numerical simulations based on Eq. (S-2), we find that three factors affect such processes. The first one of the most obvious is the saturation intensity $I_{S}$. Since, the gain rate is the constant $\kappa_{2,0}$ in the limit $I_{S}\rightarrow\infty$, it is better to have an $I_{S}$ as higher as possible, as seen from Fig. S-1(a). Due to a lower $I_{S}$, the system will evolve to a state obviously deviating from $|\langle\hat{a}_{1}\rangle|=0$. However, this defect can be alleviated by the two other factors that are relevant to the dynamical evolution of the system. In Fig. S-1(b), we fix the gain saturation intensity $I_{S}$ together with other parameter, but the initial rate $\kappa_{2,0}$ can be varied. A smaller $\kappa_{2,0}$ is beneficial to realizing a state close to the dark mode with $|\langle\hat{a}_{1}\rangle|=0$. Increasing the inter-cavity coupling $J_{1}$ between cavity $1$ and cavity $2$ can also help the realization of the dark mode; see the comparison in Fig. S-1(c) showing that a larger $J_{1}$ makes the system evolve to a more suppressed $|\langle\hat{a}_{1}\rangle|$. A better dark mode in approximation can be approached with the suitable parameters. For example, with this set of parameters, $\kappa_{2,0}=10^{-3}\,\mathrm{MHz}$, $I_{S}=10^{8}$, $J_{1}=10\,\mathrm{MHz}$, $\kappa_{1}=2\kappa_{ex}=2\,\mathrm{MHz}$, and $\kappa_{3}=1\,\mathrm{MHz}$, the system under the above-mentioned conditions and a resonantly input field with $\varepsilon_{p}=10^{4}\,\sqrt{\mathrm{MHz}}$ will have the finally evolved photon number $\langle\hat{a}_{1}^{\dagger}\hat{a}_{1}\rangle=7.69\times 10^{-6}$ in cavity $1$.
@@
An interesting point is that the exact dark mode under arbitrary gain saturation can be also realized by adjusting the coupling $J_{2}$. With the computation platforms such as Wolfram Mathematica, we can obtain the steady solution to Eq. (S-2) (after setting the left hand sides of the mean-field version of the equations to be zero) in a highly complex form of continued fraction. The coupling strength $J_{2}$ at the point of $|\langle\hat{a}_{1}\rangle|=0$ can be thus numerically found. For example, given a setup with $\kappa_{2,0}=0.01\,\mathrm{MHz}$, $I_{S}=10^{8}$, $J_{1}=10\,\mathrm{MHz}$, $\kappa_{3}=1\,\mathrm{MHz}$, and $\kappa_{1}=2\kappa_{ex}=2\,\mathrm{MHz}$, the coupling rate $J_{2}$ between cavity $2$ and $3$ is numerically found at $J_{2}=0.099\,\mathrm{MHz}$ to achieve the exact $|\langle\hat{a}_{1}\rangle|=0$ under a resonantly input field with $\varepsilon_{p}=10^{4}\,\sqrt{\mathrm{MHz}}$. This coupling strength is not far away from the value $J_{2}=\sqrt{\kappa_{2,0}\kappa_{3}}$.
@@
A.2 B. Induced Transparency with Different Structures
Without loss of generality, we here consider the resonators with the same resonance frequency $\omega_{0}$, which are coupled one after another to form a string. The induced transparency phenomenon starts from two coupled passive resonators with their damping rates $\kappa_{1}$ and $\kappa_{2}$, respectively. From the corresponding dynamics, one finds the steady intracavity field amplitude in cavity $1$ as (the notation is consistent with our form of the dynamical equations in the main text):
$A_{1}=\dfrac{\sqrt{2\kappa_{ex}}\varepsilon_{p}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{\kappa_{2}-ix}},\quad$ (S-4)
where $x=\omega_{p}-\omega_{0}$. In terms of the continued fraction, this form exactly corresponds to that of the excited level population in a three-level atomic medium for atomic EIT [22], and leads to the transmission rate $T=C^{2}/(C+1)^{2}$ at the center of the transparency window, where the cooperativity parameter is given as $C=J_{1}^{2}/(\kappa_{1}\kappa_{2})$ (reformulated according to our form of dynamical equations). Like many other types of induced transparency due to mode coupling, its level of transparency continuously improves with the increased inter-cavity coupling $J_{1}$, analogous to the EIT of three-level atomic medium where the Rabi frequency of a much stronger coupling field plays the same role as that of $J_{1}$ here, but it can never reach a perfect intracavity field cancellation to $A_{1}=0$, due to the fact $\kappa_{1},\kappa_{2}\neq 0$. If the second resonator is replaced by an active one with a constant effective gain rate $\kappa_{2}$, like in PT symmetric photonics [48, 52], the concerned intracavity field amplitude will become
$A_{1}=\dfrac{\sqrt{2\kappa_{ex}}\varepsilon_{p}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{-\kappa_{2}-ix}},\quad$ (S-5)
where the parameter $\kappa_{2}$ changes the sign. In this case, there only exists one trivial possibility of complete transparency at $x=0$, when the gain rate $g_{2}$ of the second resonator gets equal to its intrinsic damping rate $\gamma_{2}$ such that $\kappa_{2}=g_{2}-\gamma_{2}=0$.
@@
$A_{1}=\dfrac{\sqrt{2\kappa_{ex}}\varepsilon_{p}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{\kappa_{2}-ix+\dfrac{J_{2}^{2}}{\kappa_{3}-ix}}},\quad$ (S-6)
from Eq. (S-4) in terms of the continued fraction. This structure has the similar properties to those of two coupled passive resonators. However, if cavity $2$ is replaced by an active one such that $\kappa_{2}$ changes the sign to $\kappa_{2}$, i.e.,
$A_{1}=\dfrac{\sqrt{2\kappa_{ex}}\varepsilon_{p}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{-\kappa_{2}-ix+\dfrac{J_{2}^{2}}{\kappa_{3}-ix}}},\quad$ (S-7)
the subfraction below the factor $J_{1}^{2}$ in the above expression can totally vanish under the conditions $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ and $x=0$. It realizes a special dark mode with $A_{1}=0$, which will fundamentally advance instead of quantatively improving the induced transparency. In contrast to the structures of coupled passive resonators, moreover, the induced transparency with the 100100% transmission exists irrespective of the coupling strength $J_{1}$ to the first resonator.
@@
Then, the concerned intracavity field amplitude will be generalized to
$A_{1}=\dfrac{\sqrt{2\kappa_{ex}}\varepsilon_{p}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{-\kappa_{2}-ix+\dfrac{J_{2}^{2}}{\kappa_{3}-ix+\dfrac{J_{3}^{2}}{\kappa_{4}-ix}}}},\quad$ (S-9)
in terms of the corresponding form of a continued fraction. A complete transparency can be thus realized after meeting the conditions,
@@
such that the subfraction below the factor $J_{1}^{2}$ in Eq. (S-9) vanishes. Hence, the occurrence of a dark mode in cavity $1$ is completely determined by the parameters of the rest resonators—the coupling strengths $J_{2}$ and $J_{3}$ for the two pairs of the other resonators, together with the damping rates $\kappa_{3}$, $\kappa_{4}$ and gain rate $\kappa_{2}$ of them. Different from the possible field cancellation to $A_{1}=0$ due to two active-passive-coupled resonators, the gain rate $\kappa_{2}$ in this structure can have arbitrary value, as long as it locates on the submanifold determined by Eq. (S-10), which is possible by means of the adjustment of the other parameters. The generalization to even more resonators in a string is straightforward. Even if the structure is further extended beyond a string of coupled resonators, it will be still feasible to have the modified forms of continued fraction, similar to a practice in the atomic EIT involving multiple energy levels [9]. These continued fractions give us the more complex submanifolds, on which the complete transparency through the structures can be realized while the parameter spaces grow to the higher and higher dimensions.
A.3 C. Cavity Dark Mode and Effective Hamiltonian
@@
in the reference frame rotating at the pump frequency $\omega_{p}$. Like a laser beam successively split by a group of beamsplitters, the cavity field state due to the action of the above Hamiltonian is a product of coherent states $\left|\alpha_{1},\alpha_{2},\alpha_{3}\right\rangle=\left|A_{1},A_{2},A_{3}\right\rangle$, with $\hat{a}_{i}\left|A_{i}\right\rangle=A_{i}\left|A_{i}\right\rangle$ for $i=1,2,3$. Under the condition $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$, the evolved intracavity fields are the following:
@@
$A_{3}$ $=$ $-\dfrac{\sqrt{\kappa_{2}\kappa_{3}}}{J_{1}\kappa_{3}}\times\sqrt{2\kappa_{ex}}\varepsilon_{p},$ (S-12)
for three cavities with the same resonance frequency $\omega_{1}=\omega_{2}=\omega_{3}=\omega_{0}$, which are under the pump of the detuning $x=\omega_{p}-\omega_{0}=0$. The effective Hamiltonian will be thus reduced to
@@
Confined to the state $\left|0,A_{2}\right\rangle$ of the first two cavities, only the two terms, $J_{1}\hat{a}_{1}^{\dagger}\hat{a}_{2}$ and $i\sqrt{2\kappa_{ex}}\hat{a}_{1}^{\dagger}\varepsilon_{p}$, in Eq. (S-14) can affect the photon number in cavity $1$. However, their joint action exactly cancel the contribution of each other according to the solution in Eq. (S-12), implying a perfect destructive interference of the field coupled back from cavity $2$ with the input from the pump. Then, we define two Hamiltonians $\hat{H}_{1}=i\sqrt{2\kappa_{ex}}\varepsilon_{p}(\hat{a}_{1}^{\dagger}-\hat{a}_{1})$ and $\hat{H}_{2}=-J_{1}(A_{2}\hat{a}_{1}^{\dagger}+A_{2}^{\ast}\hat{a}_{1})$ of the equivalent effects for these two sources of interference. The joint action, $\langle 1|\hat{H}_{1}+\hat{H}_{2}|0\rangle=0$, forbids the excitation of cavity $1$, when the setup is under the conditions $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ and $x=0$. This result is irrespective of the parameters $J_{1}$ and $\sqrt{2\kappa_{ex}}\varepsilon_{p}$ that are directly relevant to the field in cavity $1$.
@@
A.4 D. Dynamical Stability and Experimental Feasibility
@@
The above-mentioned feasibility is relevant to the dynamical stability of the system with a constant effective gain rate $\kappa_{2}$ ($I_{S}\rightarrow\infty$), and it is determined by the eigenvalues of matrix M defined in Eq. (3) of the main text. The situations of $\kappa_{2}<\kappa_{3}$ have the totally different stable regimes from those of $\kappa_{2}>\kappa_{3}$. To the former, there are the following results for the resonant input fields:
@@
These stable regimes are exemplified in Fig. S-2(a). Here we assume that only the two coupling strengths $J_{1}$ and $J_{2}$ are tunable by the relative distances between the resonators, with the practical examples for microspheres [15] and microtoroid resonators [67] (the range of tuning $J_{1}$ and $J_{2}$ of this type of microresonators can be over $1$ GHz), while all other system parameters are fixed after the resonators are fabricated. Fig. 4 in the main text presents a process of tuning the coupling $J_{2}$ while the value of $J_{1}$ is fixed on the left side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$ in Fig. S-2(a). The coupling $J_{2}$ can be continuously tuned downwards across the horizontal line of $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ in Fig. S-2(a), which realizes a complete cancellation of the field in cavity $1$, outputting the field from lower than the $100\%$ transmission to higher than the $100\%$ transmission. After the coupling $J_{2}$ touches the lower boundary, the system will enter the regime of dynamical instability where the optical gain saturation will play a significant role so that the effective gain rate $\kappa_{2}$ cannot be treated as a constant. On the right side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$, the output field intensity at the transparency window center will tend to infinitely high as the coupling $J_{2}$ is approaching to zero, but the associate transparency window size will gradually vanish at the same time.
@@
If the built-in effective gain rate $\kappa_{2}$ is higher than the damping rate $\kappa_{3}$ in a neighboring resonator, the system will have to reduce the regimes of stability. In this case, the system stability after an increased gain rate $\kappa_{2}$ should be preserved by a correspondingly enhanced inter-cavity coupling $J_{1}$; see Fig. S-2(b). For the example in Fig. S-2(b), the coupling $J_{1}$ must be increased to $3.017$ MHz when the effective gain rate $\kappa_{2}$ reaches $1$ MHz (larger than the used $\kappa_{3}=0.1$ MHz). Because the situations of $\kappa_{2}>\kappa_{3}$ are less important, we will not elaborate on the distribution of associated dynamical regimes.

### v2

@@ Abstract
In summary, we have presented an interference mechanism of genuinely reducing the susceptibility $\varepsilon_{T}$ of the structure in Fig. 1(a), which is exactly solvable, to zero under the implementable conditions, so that the realized 100% transmission exists for any feasible power of transmitted fields. Apart from scanning the input frequency, we only require the tuning of two inter-cavity couplings to see its manifestation: setting one inter-cavity coupling at $J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ and independently adjusting the other coupling $J_{1}$ (possibly together with the external coupling $\kappa_{ex}$) for a suitable transparency window size. Rather than being used to cancel the transmission loss, the optical gain creates the point of perfect interference in the first resonator, together with two other parameters $J_{2}$ and $\kappa_{3}$ that are also totally independent of the properties of cavity $1$, and can modify the dispersion of a transmitted field at the same time. The significance of the scenario lies in two aspects. First, to the research of induced transparency based on optical resonator [22, 23], it proves that the ultimate goals of complete transparency and free control on the output/input ratio can be realized without complex designs or particular nonlinear elements. Second, it will non-trivially extend the scope of non-Hermitian photonics based on optical gain. Optical gain in one of two coupled optical resonators provides a realization of PT symmetry [47, 48, 49, 50, 51, 52, 60, 69], and the study of the associated exceptional points has been developed into a direction of intensive research [70, 71, 72]. Coupling one more resonator will have an even more non-trivial scenario involving a special dark mode with its rich properties, and it presents a rare example of something unusual simply built on the common elements like active, passive resonators and their linear couplings.
@@ Appendix A Appendix
where the noise operators satisfy the relations $\langle\hat{\xi}(t)\hat{\xi}^{\dagger}(t^{\prime})\rangle=\delta(t-t^{\prime})$ and $\langle\hat{\zeta}^{\dagger}(t)\hat{\zeta}(t^{\prime})\rangle=0$. In some other literature the damping terms take the form $\kappa/2$ while the corresponding noise terms have the factor $\sqrt{\kappa}$. Particularly, we consider the gain saturation effect such that the net gain rate with time takes the form
@@
In some applications, gain saturation effect leads to optical isolation of coupled optical resonators [53, 56, 60], modifies optomechanical oscillations of hybrid systems [55, 57], and can also influence macroscopic quantum properties of coupled resonators [54]. In our concerned system of three coupled optical resonators, the time-dependent net optical gain rate can affect the intracavity fields. The indicator for the system to evolve to a desired dark mode is whether the field amplitude in the first cavity will finally tend to $|\langle\hat{a}_{1}\rangle|=0$, where $\langle\hat{a}_{1}\rangle=A_{1}e^{-i\omega_{p}t}$. Without a loss of the described features, we approximate with $\gamma_{2}=0$ in our model calculations because this loss can be compensated by a higher $\kappa_{2,0}$, so that the net gain rate keeps to be $\kappa_{2}(t)>0$. We let the system evolve under two conditions: (1) the resonant input for three identical resonance cavity frequencies; (2) $J_{2}=\sqrt{\kappa_{2,0}\kappa_{3}}$ for the chosen system parameters. For a constant $\kappa_{2}=\kappa_{2,0}$, there will emerge an exact $|\langle\hat{a}_{1}\rangle|=0$ under these conditions, as we illustrate in the main text, but this result cannot be guaranteed with a variable $\kappa_{2}(t)$ in Eq. (S-3). Through the numerical simulations based on Eq. (S-2), we find that three factors affect such processes. The first one of the most obvious is the saturation intensity $I_{S}$. Since the gain rate is a constant $\kappa_{2,0}$ in the limit $I_{S}\rightarrow\infty$, it is better to have an $I_{S}$ as higher as possible, as seen from Fig. S-1(a). Due to a lower $I_{S}$, the system will evolve to a state obviously deviating from $|\langle\hat{a}_{1}\rangle|=0$. However, this defect can be alleviated by the two other factors that are relevant to the dynamical evolution of the system. In Fig. S-1(b), we fix the gain saturation intensity $I_{S}$ together with other parameter, but the initial rate $\kappa_{2,0}$ can be varied. A smaller $\kappa_{2,0}$ is beneficial to realizing a state close to the dark mode with $|\langle\hat{a}_{1}\rangle|=0$. Increasing the inter-cavity coupling $J_{1}$ between cavity $1$ and cavity $2$ can also help the realization of a better dark mode; see the comparison in Fig. S-1(c) showing that a larger $J_{1}$ makes the system evolve to a more suppressed $|\langle\hat{a}_{1}\rangle|$. A better dark mode in approximation can be approached with the suitable parameters. For example, with this set of parameters, $\kappa_{2,0}=10^{-3}\,\mathrm{MHz}$, $I_{S}=10^{8}$, $J_{1}=10\,\mathrm{MHz}$, $\kappa_{1}=2\kappa_{ex}=2\,\mathrm{MHz}$, and $\kappa_{3}=1\,\mathrm{MHz}$, the system under the above-mentioned conditions and a resonantly input field with $\varepsilon_{p}=10^{4}\,\sqrt{\mathrm{MHz}}$ will have the finally evolved photon number $\langle\hat{a}_{1}^{\dagger}\hat{a}_{1}\rangle=7.69\times 10^{-6}$ in cavity $1$.
@@
An interesting point is that the exact dark mode under an arbitrary gain saturation can be also realized by adjusting the coupling $J_{2}$. With the computation platforms such as Wolfram Mathematica, we can obtain the steady solution to Eq. (S-2) (after setting the left hand sides of the mean-field version of the equations to be zero) in a highly complex form of continued fraction. The coupling strength $J_{2}$ at the point of $|\langle\hat{a}_{1}\rangle|=0$ can be thus numerically found. For example, given a setup with $\kappa_{2,0}=0.01\,\mathrm{MHz}$, $I_{S}=10^{8}$, $J_{1}=10\,\mathrm{MHz}$, $\kappa_{3}=1\,\mathrm{MHz}$, and $\kappa_{1}=2\kappa_{ex}=2\,\mathrm{MHz}$, the coupling rate $J_{2}$ between cavity $2$ and $3$ is numerically found at $J_{2}=0.099\,\mathrm{MHz}$ to achieve the exact $|\langle\hat{a}_{1}\rangle|=0$ under a resonantly inputting field with $\varepsilon_{p}=10^{4}\,\sqrt{\mathrm{MHz}}$. This coupling strength is not far away from the value $J_{2}=\sqrt{\kappa_{2,0}\kappa_{3}}$.
@@
Appendix B B. Induced Transparency with Different Structures
Without loss of generality, we here consider that the resonators, which are coupled one after another to form a string, have the same resonance frequency $\omega_{0}$ and the net gain rate $\kappa_{2}$ used in some of the structures is a constant. The induced transparency phenomenon starts from two coupled passive resonators with their damping rates $\kappa_{1}$ and $\kappa_{2}$, respectively. From the corresponding dynamics, one finds the susceptibility of the structure as
$\varepsilon_{T}=\frac{\sqrt{2\kappa_{ex}}A_{1}}{\varepsilon_{p}}=\dfrac{2\kappa_{ex}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{\kappa_{2}-ix}},\quad$ (S-4)
where $x=\omega_{p}-\omega_{0}$ and the notation is consistent with our form of the dynamical equations in the main text. The above form exactly corresponds to that of a three-level atomic medium for atomic EIT [22], and it leads to the transmission rate $T=C^{2}/(C+1)^{2}$ at the center of transparency window, where the cooperativity parameter is given as $C=J_{1}^{2}/(\kappa_{1}\kappa_{2})$ (reformulated according to our form of dynamical equations). Its level of transparency continuously improves with the increased inter-cavity coupling $J_{1}$, analogous to the EIT of three-level atomic medium where the Rabi frequency of a much stronger coupling field plays the same role as that of $J_{1}$ here, but it can never reach a perfect intracavity field cancellation to $\varepsilon_{T}=0$, due to the fact $\kappa_{1},\kappa_{2}\neq 0$. The complete transparency (100100% transmission) with this structure can be only asymptotically approached with nearly lossless resonators or huge inter-cavity coupling strength. If the second resonator is replaced by an active one with a constant net gain rate $\kappa_{2}$, like in PT symmetric photonics [48, 52], the susceptibility will become
$\varepsilon_{T}=\dfrac{2\kappa_{ex}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{-\kappa_{2}-ix}},\quad$ (S-5)
where the parameter $\kappa_{2}$ changes the sign. In this case, there only exists one trivial possibility of complete transparency at $x=0$, when the gain rate $g_{2}$ of the second resonator gets equal to its intrinsic damping rate $\gamma_{2}$ such that $\kappa_{2}=g_{2}-\gamma_{2}=0$. It should be noted that the dispersion properties reflected by the imaginary part of the susceptibility cannot be modified by the added gain, so that the induced transparency and field amplification work separately in this configuration.
@@
$\varepsilon_{T}=\dfrac{2\kappa_{ex}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{\kappa_{2}-ix+\dfrac{J_{2}^{2}}{\kappa_{3}-ix}}}\quad$ (S-6)
from Eq. (S-4). This structure has the similar properties to those of two coupled passive resonators. However, if cavity $2$ is replaced by an active one such that $\kappa_{2}$ changes the sign to $\kappa_{2}$, i.e.,
$\varepsilon_{T}=\dfrac{2\kappa_{ex}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{-\kappa_{2}-ix+\dfrac{J_{2}^{2}}{\kappa_{3}-ix}}},\quad$ (S-7)
the part below the factor $J_{1}^{2}$ in the above expression can totally vanish under the conditions $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ and $x=0$. It realizes a special dark mode with $\langle\hat{a}_{1}\rangle=0$, which fundamentally advances the induced transparency with coupled optical resonators. In contrast to the structures of coupled passive resonators, moreover, the induced transparency with the 100100% transmission exists irrespective of the coupling strength $J_{1}$ to the first resonator.
@@
Then, the susceptibility will be further generalized to
$\varepsilon_{T}=\dfrac{2\kappa_{ex}}{\kappa_{1}-ix+\dfrac{J_{1}^{2}}{-\kappa_{2}-ix+\dfrac{J_{2}^{2}}{\kappa_{3}-ix+\dfrac{J_{3}^{2}}{\kappa_{4}-ix}}}}.\quad$ (S-9)
A complete transparency can be thus realized after meeting the conditions:
@@
such that the part below the factor $J_{1}^{2}$ in Eq. (S-9) vanishes. Hence, the occurrence of a dark mode in cavity $1$ is completely determined by the parameters of the rest resonators—the coupling strengths $J_{2}$ and $J_{3}$ for the two pairs of the other resonators, together with the damping rates $\kappa_{3}$, $\kappa_{4}$ and gain rate $\kappa_{2}$ of them. Different from the possible field cancellation to $\langle\hat{a}_{1}\rangle=0$ due to two active-passive-coupled resonators, the net gain rate $\kappa_{2}$ in this structure can have arbitrary value, as long as it locates on the submanifold determined by Eq. (S-10), which is possible by means of the adjustment of the other parameters. The generalization to even more resonators in a string is straightforward. Even if the structure is further extended beyond a string of coupled resonators, it will be still feasible to have the modified forms of continued fraction, similar to a practice in the atomic EIT involving multiple energy levels [9]. These continued fractions give us the more complex submanifolds, on which the complete transparency through the structures can be realized while the parameter spaces grow to the higher and higher dimensions.
Appendix C C. Cavity Dark mode and Effective Hamiltonian
@@
in the reference frame rotating at the input field frequency $\omega_{p}$. Like a laser beam successively split by a group of beamsplitters, the cavity field state due to the action of the above Hamiltonian is a product of coherent states $\left|\alpha_{1},\alpha_{2},\alpha_{3}\right\rangle=\left|A_{1},A_{2},A_{3}\right\rangle$, with $\hat{a}_{i}\left|A_{i}\right\rangle=A_{i}\left|A_{i}\right\rangle$ for $i=1,2,3$. Under the condition $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$, the evolved intracavity fields are the following:
@@
$A_{3}$ $=$ $-\sqrt{\frac{\kappa_{2}}{\kappa_{3}}}\frac{1}{J_{1}}\times\sqrt{2\kappa_{ex}}\varepsilon_{p},$ (S-12)
for three cavities with the same resonance frequency $\omega_{1}=\omega_{2}=\omega_{3}=\omega_{0}$, which are under the driving field of the detuning $x=\omega_{p}-\omega_{0}=0$. The effective Hamiltonian will be thus reduced to
@@
Confined to the state $\left|0,A_{2}\right\rangle$ of the first two cavities, only the two terms, $J_{1}\hat{a}_{1}^{\dagger}\hat{a}_{2}$ and $i\sqrt{2\kappa_{ex}}\hat{a}_{1}^{\dagger}\varepsilon_{p}$, in Eq. (S-14) can affect the photon number in cavity $1$. However, their joint action exactly cancel the contribution of each other according to the solution in Eq. (S-12), implying a perfect destructive interference of the field coupled back from cavity $2$ with the inputted field. Then, we define two Hamiltonians $\hat{H}_{1}=i\sqrt{2\kappa_{ex}}\varepsilon_{p}(\hat{a}_{1}^{\dagger}-\hat{a}_{1})$ and $\hat{H}_{2}=-J_{1}(A_{2}\hat{a}_{1}^{\dagger}+A_{2}^{\ast}\hat{a}_{1})$ of the equivalent effects for these two sources of interference. The joint action, $\langle 1|\hat{H}_{1}+\hat{H}_{2}|0\rangle=0$, forbids the excitation of cavity $1$, when the setup is under the conditions $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ and $x=0$. This result is irrespective of the parameters $J_{1}$ and $\sqrt{2\kappa_{ex}}\varepsilon_{p}$ that are directly relevant to the field in cavity $1$.
@@
C.1 D. Dynamical Stability and Experimental Feasibility
@@
The above-mentioned feasibility is relevant to the dynamical stability of the system with a constant net gain rate $\kappa_{2}$ ($I_{S}\rightarrow\infty$), and it is determined by the eigenvalues of matrix M defined in Eq. (3) of the main text. The situations of $\kappa_{2}<\kappa_{3}$ have the totally different stable regimes from those of $\kappa_{2}>\kappa_{3}$. To the former, there are the following results for the resonant input fields:
@@
These stable regimes are exemplified in Fig. S-2(a). Here we assume that only the two coupling strengths $J_{1}$ and $J_{2}$ are tunable by the relative distances between the resonators, with the practical examples for microspheres [15] and microtoroid resonators [67] (the range of tuning $J_{1}$ and $J_{2}$ of this type of microresonators can be over $1$ GHz), while all other system parameters are fixed after the resonators are fabricated. Fig. 4 in the main text presents a process of tuning the coupling $J_{2}$ while the value of $J_{1}$ is fixed on the left side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$ in Fig. S-2(a). The coupling $J_{2}$ can be continuously tuned downwards across the horizontal line of $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ in Fig. S-2(a), on which there is a complete cancellation of the field in cavity $1$, outputting the field from lower than the $100\%$ transmission to higher than the $100\%$ transmission. After the coupling $J_{2}$ touches the lower boundary, the system will enter the regime of dynamical instability, where the optical gain saturation will play a significant role so that the finally stabilized gain rate will be much lower than the original $\kappa_{2,0}$. On the right side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$, the output field intensity at the transparency window center will tend to infinitely high as the coupling $J_{2}$ is approaching to zero, but the associate transparency window will be vanishing at the same time.
@@
If the built-in net gain rate $\kappa_{2}$ is higher than the damping rate $\kappa_{3}$ in a neighboring resonator, the system will have to reduce the regimes of stability. In this case, the system stability after an increased gain rate $\kappa_{2}$ should be preserved by a correspondingly enhanced inter-cavity coupling $J_{1}$; see Fig. S-2(b). For the example in Fig. S-2(b), the coupling $J_{1}$ must be increased to $3.017$ MHz when the net gain rate $\kappa_{2}$ reaches $1$ MHz (larger than the used $\kappa_{3}=0.1$ MHz). Because the situations of $\kappa_{2}>\kappa_{3}$ are less important, we will not elaborate on the distribution of associated dynamical regimes.